I have three fraction questions that I need help with because I'm not good with fractions at all the question are,solve the equation for b, 11over12+b=5over6 and I don't know where to start.The next one, solve the equation for f, f-(-7over10)=-4over5.And the last one is, if c=5over6 and d=1 over 4,evaluate 8c over-d and that is the last question thank you for any help I can get.
Answer by stanbon(75887)

I have three fraction questions that I need help with because I'm not good with fractions at all the question are,solve the equation for b, 11over12+b=5over6 and I don't know where to start.
------------
11/12 + b = 5/6
Multiply thru by 12 to get:
11 + 12b = 10
12b = -1
b = -1/12
=====================
The next one, solve the equation for f, f-(-7over10)=-4over5.
f-(-7/10) = (-4/5)
MUltiply thru by 10 to get:
10f - (-7) = -8
10f = -15
f = -3/2
===================
And the last one is, if c=5over6 and d=1 over 4,evaluate 8c over-d and that is the last question thank you for any help I can get.
c=5/6 and d=1/4,
Then (8c)/(-d)
=8(5/6)/(-1/4)
=(20/3)/(-1/4)
Invert the denominator and multiply to get:
= - 80/3
==============
Cheers,
Stan H.
-------------